OpenStudy (anonymous):

X is a normally distributed random variable with a mean of 11.0 and a standard deviation of 3.50. Find the value x such that P(X < x) is equal to 0.86

OpenStudy (anonymous):

Let \(Y\) denote the standardized normal distribution, which has \(Z=\dfrac{X-\mu}{\sigma}\). So, you have \[P(X<x)=0.86=P(\mu+\sigma Z<x)=P(11+3.5^2Z<x)\] Using a table for \(z\)-values, you can find that the appropriate \(z\) that gives you 0.86, so you would just have to solve for \(x\): \[11+3.5^2z=x\]

OpenStudy (anonymous):

thank yo u